# Self-Supervised Speed of Sound Recovery for Aberration-Corrected Photoacoustic Computed Tomography

### v3

@@ Abstract
Photoacoustic computed tomography (PACT) is a non-invasive imaging modality, similar to ultrasound, with wide-ranging medical applications. Conventional PACT images are degraded by wavefront distortion caused by the heterogeneous speed of sound (SOS) in tissue. Accounting for these effects can improve image quality and provide medically useful information, but measuring the SOS directly is burdensome and the existing joint reconstruction method is computationally expensive. Traditional supervised learning techniques are currently inaccessible in this data-starved domain. In this work, we introduce an efficient, self-supervised joint reconstruction method that recovers SOS and high-quality images using a differentiable physics model to solve the semi-blind inverse problem. The SOS, parametrized by either a pixel grid or a neural field (NF), is updated directly by backpropagation. Our method removes SOS aberrations more accurately and 35x faster than the current SOTA. We demonstrate the success of our method quantitatively in simulation and qualitatively on experimentally-collected and in-vivo data.
Figure 1: Self-Supervised Speed of Sound Recovery Corrects PACT Aberrations. (a) Backprojection converts photoacoustic measurements into images (purple), which suffer aberrations caused by unaccounted-for variation in the speed of sound (SOS) of tissue. Our method produces high-quality images through a self-supervised recovery of SOS. We begin with a stack of aberrated reconstructions, created with a varying delay parameter (green), roughly analogous to a focal stack. Each image patch is recovered with a multi-channel deconvolution. Our forward model (dashed red) uses a trainable SOS representation to estimate the point spread functions (PSFs) at each location (white “x” for example shown). The physics-based forward model and multichannel inversion are fully differentiable, enabling test-time training without any external training data, which is currently limited for this imaging modality. (b) We benchmark our method using both a neural field (NF) and pixel grid (PG) for SOS representation. Compared to the best existing methods, both with and without SOS recovery, we provide state-of-the-art image quality, and recover SOS with an order-of-magnitude speed-up.
@@ 1 Introduction
Though supervised learning methods have made significant progress on image restoration in many applications, PACT is a relatively novel imaging modality and appropriate datasets (large and naturally distributed set of measurements, with realistic aberrations, accompanied by ground truth for both image and SOS values) are not available. We address these challenges with a self-supervised joint recovery method that learns SOS by backpropagation through a differentiable physical model, then uses the SOS to reconstruct a clean image. Specifically, as shown in in Fig. 1, we optimize a differentiable representation of the SOS, parameterized by a pixel grid (PG) or neural field (NF), to best match observed aberrations, which can then be removed with deconvolution. By imposing physical consistency across image locations and a delay-based “focal stack” (details in Sec. 3), this method converges quickly to produce an approximate SOS and high quality image. Our key contributions are:
@@
We benchmark the proposed method and previous methods on simulated data, showing that our method achieves state-of-the-art reconstructions in a fraction of the compute time (40 sec vs. 23 min). In qualitative results on real data, we show the method is robust to nonideal effects (e.g. transducer responses and geometric offsets) found in PACT systems, as demonstrated by its translation to in vivo data without modification.
@@ 2 Related Work
In contrast to methods that assume or measure SOS, joint reconstruction approaches simultaneously recover images and 2D SOS from PA signals. Regularization-based joint reconstruction [45, 25] alternatively reconstructs the image and SOS in an iterative framework with regularization on both distributions. However, the ill-posed nature of the problem causes numerical instability and error accumulation. The feature coupling method [3] parameterizes the SOS with several constant value regions and maximizes the similarity between reconstructed images from partial arrays. Its drawbacks are the inability to model a complicated SOS map and reliance on a good initial guess. Inspired by indirect wavefront sensing in adaptive optics (AO) [14, 9, 46], Adaptive PACT (APACT) [6] divides the image into small patches and solves for the wavefront in every patch via an exhaustive search. Of these methods, APACT produces state-of-the-art image quality, but with two major drawbacks. First, APACT assumes a simplified wavefront model that cannot account for overall scaling of the reconstructed image. Second, its exhaustive search is computationally expensive, leading to extended reconstruction times that limit its practical application. Our method addresses these weaknesses with a novel self-supervised, physics-based deep learning approach.
@@ 3.1 Image Formation and Aberration
Filtered back-projection [40] is a widely used image reconstruction method in ring array PACT systems. In practice, due to the limited bandwidth and non-flat angular response of the transducers [6], a delay-and-sum (DAS) algorithm [31, 10, 12] is often used as a reasonable approximation. A delay-and-sum image $y(\mathbf{r};d)$ varies over sample locations $\mathbf{r}^{\prime}$ as well as a user-selected extra delay parameter $d$ that functions as a “focus setting” [33], and is reconstructed according to
@@
Figure 2: Acoustic heterogeneity creates non-ideal PSFs due to wavefront distortions. We consider three SOS patterns (a): the uniform SOS assumed by standard methods, a simplified body model with two internal SOS values, and a more realistic numerical phantom that we created to mimic a body slice including representative materials ranging from air to bone. For the location marked with a white x, we illustrate (b) the discrepancy between the standard methods’ assumed wavefront and the true wavefront location arising from the SOS shown. These wavefront errors determine the PSFs (c) at the indicated location for each SOS, which lead to image aberrations for non-pointlike PSFs. A user-set delay value in the image reconstruction changes the resulting PSFs and aberrations (d), and can be used to effectively remove aberrations for simple SOS models but not realistic ones.
@@
where $\theta$ is the angle of the vector $\mathbf{r}-\mathbf{r}^{\prime}$. See Fig. 2b for example wavefront errors corresponding to the SOS maps $\mathbf{v}$ shown in panel a for the location $\mathbf{r}^{\prime}$ indicated by white crosses. It has been shown in [6] that the resulting aberrations, despite varying across the sample, can be modeled within isoplanatic image patches $y_{i}$ as convolutions with PSFs $h_{i}$ and unaberrated patches $x_{i}$:
@@ 3.2 Multi-channel Deconvolution
Fig. 3b shows the image patch recovered using single-channel deconvolution with each delay’s PSF. The most compact PSF (red) produces the best result of these, but artifacts are extreme for each individual channel. By using 16 delays for multichannel deconvolution, our method reduces ringing and includes features that are otherwise lost at patch edges, see arrows in Fig. 3c. The reduction of aberrations as more delay channels are added can be seen in Fig. 3d, which also shows a growth in compute time from 0.9 to 5.0 seconds for 4 to 32 channels.
Figure 4: Numerical simulation demonstrates accuracy and computational efficiency. (a) Our numerical phantom allows quantitative evaluation of image reconstruction under different SOS distributions in simulation. The PA signals are generated with k-wave simulation [32] using the IP and SOS shown. (b) The wavefront (sampled location marked by white “x” in a) reconstructed by our method with NF (purple) and PG (green) is much closer to the true wavefront (red) than other methods, better capturing image aberrations. (c) We compare our method with conventional DAS, Dual-SOS DAS, our multi-channel deconvolution with the true SOS, and APACT. While conventional DAS assumes a uniform SOS and results in significant image aberrations, Dual-SOS DAS provides better contrast by assuming a single-body SOS but cannot fully undo the aberrations and reconstruct the SOS. Given an accurate SOS, our multi-channel deconvolution is able to efficiently reconstruct a high-quality IP image. APACT is able to deblur the image effectively but does not address overall shrinkage (see vertical shift in green boxes) and cannot reconstruct an accurate SOS. Our method with NF solves the shrinking problem by fully characterizing the wavefront and is able to fully reconstruct the IP and SOS (see metrics) in a much shorter time than APACT. Ours with PG also produces good IP images but generates stripe-like artifacts in the SOS that originate from the path-integral nature of the wavefront computation. Providing our method with the ground truth SOS leads to only marginal improvements in image quality. See visualizations of the convergence in the supplementary videos.
@@ 3.3 Joint Reconstruction
By combining our differentiable PSF generation and deconvolution modules, shown in red in Fig. 1, we are able to backpropagate through the entire forward and inverse models to achieve consistency between the expected and observed image aberrations. This allows joint reconstruction of the unaberrated image and the SOS from PA signals with no external training data by leveraging the redundant information encoded in the DAS stack.
We benchmark two types of SOS representation: a pixel grid (PG) with the same resolution as the IP image, and a neural field (NF) representation. For the latter, we used a single-layer, 256-feature, fully-connected network with sinusoidal activations [27]. This non-pretrained, coordinate-based network maps pixel coordinates to SOS predictions. The total number of NF parameters (1027) is much smaller than the number of pixels ($\sim$200k) in the SOS map, which reduces the degrees of freedom in the inverse problem and acts as an implicit regularizer that makes the problem better-posed. See Sec. S3.3 for ablations on network size. For both representations, we reduce unnecessary computation by restricting these pixel grid or NF coordinates to a circular mask roughly the size of the sample and fix the background SOS based on measured water temperature [20]
@@
over all $N$ image patches simultaneously, where $\phi$ is the parameters in the PG or NF. This loss function imposes a physics-informed data fidelity that includes both the forward model in $\mathbf{H}_{i}$ as well as a Fourier-noise-specific $|\mathbf{k}|$ weighting as in [6]. We use TV loss to regularize the conventional pixel grid representation, while for the NF representation, we rely instead on the network’s implicit regularization and set $\lambda=0$. Training is performed only on the SOS representation parameters $\phi$ but requiring accurate aberration reconstruction across all delays and image patches leads to sufficiently accurate PSFs that the final multichannel deconvolution produces high-quality recovered IP image patches $x_{i}$. Finally, these patches are merged into a single image by applying a Gaussian window with a FWHM of 1.5 mm to the image patches [6] for smooth interpatch transitions and adding the patches back to the IP image at corresponding patch locations. We use a delay count of $M=16$ by default in this paper (see ablations in Sec. 4.4). See Sec. S1.4 for details on our image patches and stitching.
Our method is implemented with PyTorch [22] and available on GitHub (see project page). We trained for 10 epochs for NF and 30 epochs for PG with Adam optimizer [17] and learning rates of $0.1$ and $5\times 10^{-3}$ respectively on an NVIDIA RTX A6000 GPU.
@@ 4.1 Numerical Simulations
To quantitatively benchmark the performance of our method with ground truth references, we designed a set of numerical phantoms with 5 different SOS maps (see Sec. S1.1 for details), one of which is shown in Fig. 4a for visualization of the results.
@@
The wavefront errors modeled/reconstructed (sample location marked by white “x” on SOS in Fig. 4a) by different methods are shown in Fig. 4b. To use the terminology of [6], the wavefront errors can be decomposed into a Fourier series, where the components degrade the IP image differently. Conventional DAS assumes a uniform SOS thus a constant wavefront (red), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (yellow). APACT (green) includes only the zeroth and second order Fourier series terms in its wavefront model (see [6] and Sec. S2 for details), missing both the first order (the cosine component captured by dual-SOS in yellow) as well as higher order terms capturing finer detail. In comparison, our methods can accurately model lower and higher-order components in the wavefront (purple and blue), compared with the truth wavefront (black).
@@
These wavefront errors are directly reflected in image reconstruction quality in Fig. 4c. Conventional DAS, which assumes a uniform SOS of 1510.0 m/s, not only suffers from doubling artifacts (see insets) due to the ignorance of higher order terms but also causes the image to shrink (see vertical shift of the circle and boundary in green box). This shrinking is caused by the missing first-order term in the wavefront, which leads to off-center PSFs that shift the image patches toward the center of the tissue. Note also that for real samples this uniform SOS value is tuned by generating many reconstructions and requiring the user to select the best image, significantly increasing the reconstruction time in practice. Dual-SOS DAS assumes a body SOS of 1561.0 m/s (also hand-tuned with search time not included in reported time) and is able to fix the shrinking problem but artifacts remain. The image reconstructed by APACT is also shrunk (see green box and the image quality metrics) due to ignorance of the first-order term. APACT oversmooths the SOS and loses key features in the SOS, e.g., the orange liver-like region and the hole in the upper right, and is also slow and expensive due to the exhaustive search for wavefronts coefficients. Our method reconstructs a more accurate SOS, corresponding to a significant improvement in image quality over traditional methods. Image patches show fewer aberrations than all previous methods, reflected in high PSNR and SSIM scores. Our method with NF takes only 40 seconds, compared to 22 minutes for APACT. Our method with PG produces decent IP images in similar compute time but suffers from stripe-like SOS artifacts that originate from the path-integral nature of the wavefront calculation (see Eq. 3) in the straight acoustic ray model. We also show the image reconstructed by multi-channel deconvolution using the ground truth SOS as an upper limit baseline for image quality and speed of USCT-enhanced reconstruction. See Tab. 1 for quantitative evaluation of all the methods, averaged over five numerical phantoms. Supplementary videos show the convergence of the IP image and SOS for each experiment.

### v4

# Coordinate-based Speed of Sound Recovery for Aberration-Corrected Photoacoustic Computed Tomography

@@ Abstract
Photoacoustic computed tomography (PACT) is a non-invasive imaging modality, similar to ultrasound, with wide-ranging medical applications. Conventional PACT images are degraded by wavefront distortion caused by the heterogeneous speed of sound (SOS) in tissue. Accounting for these effects can improve image quality and provide medically useful information, but measuring the SOS directly is burdensome and the existing joint reconstruction method is computationally expensive. Traditional supervised learning techniques are currently inaccessible in this data-starved domain. In this work, we introduce an efficient, self-supervised joint reconstruction method that recovers SOS and high-quality images for ring array PACT systems. To solve this semi-blind inverse problem, we parametrize the SOS using either a pixel grid or a neural field (NF) and update it directly by backpropagating the gradients through a differentiable imaging forward model. Our method removes SOS aberrations more accurately and 35x faster than the current SOTA. We demonstrate the success of our method quantitatively in simulation and qualitatively on experimentally-collected and in vivo data. Our code and synthetic numerical phantoms are available on our project page: https://lukeli0425.github.io/Coord-SoS-PACT/.
Figure 1: Self-Supervised Speed of Sound Recovery Corrects PACT Aberrations. (a) Backprojection converts photoacoustic measurements into images (purple), which suffer aberrations caused by unaccounted-for variation in the speed of sound (SOS) of tissue. Our method produces high-quality images through a self-supervised recovery of SOS. We begin with a stack of aberrated reconstructions, created with a varying delay parameter (green), roughly analogous to a focal stack. Each image patch is recovered with a multi-channel deconvolution. Our forward model (dashed red) uses a trainable coordinate-based SOS representation to estimate the point spread functions (PSFs) at each location (white “x" for example shown). The physics-based forward model and multi-channel inversion are fully differentiable, enabling test-time training without any external training data, which is currently limited for this imaging modality. (b) We benchmark our method using both a neural field (NF) and pixel grid (PG) for SOS representation. Compared to the best existing methods, both with and without SOS recovery, we provide state-of-the-art image quality, and recover SOS with an order-of-magnitude speed-up.
@@ 1 Introduction
Though supervised learning methods have made significant progress on image restoration in many applications, PACT is a relatively novel imaging modality and appropriate datasets (large and naturally distributed set of measurements, with realistic aberrations, accompanied by ground truth for both image and SOS values) are not available. We consider a ring array PACT system and address these challenges with a self-supervised joint recovery method that learns SOS by backpropagation through a differentiable physical model, then uses the SOS to reconstruct a clean image. Specifically, as shown in in Fig. 1, we optimize a differentiable coordinate-based representation of the SOS, parameterized by a pixel grid (PG) or neural field (NF), to best match observed aberrations, which can then be removed with deconvolution. By imposing physical consistency across image locations and a delay-based “focal stack" (details in Sec. 3), this method converges quickly to produce an approximate SOS and high quality image. Our key contributions are:
@@
We benchmark the proposed method and previous methods on simulated data, showing that our method achieves state-of-the-art reconstructions in a fraction of the compute time (40 seconds vs. 23 minutes). In qualitative results on real data, we show the method is robust to nonideal effects (e.g. transducer responses and geometric offsets) found in PACT systems, as demonstrated by its translation to in vivo data without modification.
@@ 2 Related Work
In contrast to methods that assume or measure SOS, joint reconstruction approaches simultaneously recover images and 2D SOS from PA signals. Regularization-based joint reconstruction [44, 24] alternatively reconstructs the image and SOS in an iterative framework with regularization on both distributions. However, the ill-posed nature of the problem causes numerical instability and error accumulation. The feature coupling method [3] parameterizes the SOS with several constant value regions and maximizes the similarity between reconstructed images from partial arrays, but is unable to model a complicated SOS map and relies heavily on a good initial guess. Inspired by indirect wavefront sensing in adaptive optics (AO) [13, 8, 45], Adaptive PACT (APACT) [5] divides the image into small patches and solves for the wavefront in every patch via an exhaustive search. Of these methods, APACT produces state-of-the-art image quality, but with two major drawbacks. First, APACT assumes a simplified wavefront model that cannot account for overall scaling of the reconstructed image. Second, its exhaustive search is computationally expensive, leading to extended reconstruction times that limit its practical application. Our method addresses these weaknesses with a novel physics-based, self-supervised learning approach.
@@ 3.1 Image Formation and Aberration
Filtered back-projection [39] is a widely used image reconstruction method in ring array PACT systems. In practice, due to the limited bandwidth and non-flat angular response of the transducers [5], a delay-and-sum (DAS) algorithm [30, 9, 11] is often used as a reasonable approximation. A delay-and-sum image $y(\mathbf{r};d)$ varies over sample locations $\mathbf{r}^{\prime}$ as well as a user-selected extra delay parameter $d$ that functions as a “focus setting" [32], and is reconstructed according to
@@
Figure 2: Acoustic heterogeneity creates non-ideal PSFs due to wavefront distortions. We consider three SOS patterns (a): the uniform SOS assumed by standard methods, a simplified body model with two internal SOS values, and a more realistic numerical phantom that we created to mimic a body slice including representative materials ranging from air to bone. For the location marked with a white “x", we illustrate (b) the discrepancy between the standard methods’ assumed wavefront and the true wavefront location arising from the SOS shown. These wavefront errors determine the PSFs (c) at the indicated location for each SOS, which lead to image aberrations for non-pointlike PSFs. A user-set delay value in the image reconstruction changes the resulting PSFs and aberrations (d), and can be used to effectively remove aberrations for simple SOS models but not realistic ones.
@@
where $\theta$ is the angle of the vector $\mathbf{r}-\mathbf{r}^{\prime}$. See Fig. 2b for example wavefront errors corresponding to the SOS maps $\mathbf{v}$ shown in Fig. 2a for the location $\mathbf{r}^{\prime}$ indicated by a white “x". It has been shown in [5] that the resulting aberrations, despite varying across the sample, can be modeled within isoplanatic image patches $y_{i}$ as convolutions with PSFs $h_{i}$ and unaberrated patches $x_{i}$:
@@ 3.2 Multi-channel Deconvolution
Fig. 3b shows the image patch recovered using single-channel deconvolution with each delay’s PSF. The most compact PSF (red) produces the best result of these, but artifacts are extreme for each individual channel. By using 16 delays for multi-channel deconvolution, our method reduces ringing and includes features that are otherwise lost at patch edges, see arrows in Fig. 3c. The reduction of aberrations as more delay channels are added can be seen in Fig. 3d, which also shows a growth in compute time from 0.9 to 5.0 seconds for 4 to 32 channels.
Figure 4: Numerical simulation demonstrates accuracy and computational efficiency. (a) Our numerical phantom allows quantitative evaluation of image reconstruction under different SOS distributions in simulation. The PA signals are generated with k-Wave simulation [31] using the IP and SOS shown. (b) The wavefront (sampled location marked by white “x" in a) reconstructed by our method with NF (purple) and PG (green) is much closer to the true wavefront (red) than other methods, better capturing image aberrations. (c) We compare our method with conventional DAS, Dual-SOS DAS, our multi-channel deconvolution with the true SOS, and APACT. While conventional DAS assumes a uniform SOS and results in significant image aberrations, Dual-SOS DAS provides better contrast by assuming a single-body SOS but cannot fully undo the aberrations and reconstruct the SOS. Given an accurate SOS, our multi-channel deconvolution is able to efficiently reconstruct a high-quality IP image. APACT is able to deblur the image effectively but does not address overall shrinkage (see vertical shift in green boxes) and cannot reconstruct an accurate SOS. Our method with NF solves the shrinking problem by fully characterizing the wavefront and is able to fully reconstruct the IP and SOS (see metrics) in a much shorter time than APACT. Ours with PG also produces good IP images but generates stripe-like artifacts in the SOS that originate from the path-integral nature of the wavefront computation. Providing our method with the ground truth SOS leads to only marginal improvements in image quality. See visualizations of the convergence in the supplementary videos.
@@ 3.3 Joint Reconstruction
By combining our differentiable PSF generation and deconvolution modules, shown in red in Fig. 1, we are able to backpropagate through the entire forward and inverse models to achieve consistency between the expected and observed image aberrations.
Note in Eq. 3 that our wavefronts are calculated from the entire SOS map via path integral, therefore each incorrect SOS pixel will contribute to an incorrect PSF at every image patch through the wavefront in some direction. While the PSFs will be included in both the forward and inverse model, the overdetermined nature of Eq. 6 brings the PSFs into consistency with each other as iterations progress, enabling joint reconstruction of the unaberrated image and the SOS from PA signals alone.
We benchmark two types of SOS representation: a pixel grid (PG) with the same resolution as the IP image, and a neural field (NF) representation. For the latter, we used a single-layer, 256-feature, fully-connected network with sinusoidal activations [26]. This non-pretrained, coordinate-based network maps pixel coordinates to SOS predictions. The total number of NF parameters (1027) is much smaller than the number of pixels ($\sim$200k) in the SOS map, which reduces the degrees of freedom in the inverse problem and acts as an implicit regularizer that makes the problem better-posed. See Sec. S3.3 for ablations on network size. For both representations, we reduce unnecessary computation by restricting these pixel grid or NF coordinates to a circular mask roughly the size of the sample and fix the background SOS based on measured water temperature [19].
@@
over all $N$ image patches simultaneously, where $\phi$ is the parameters in the PG or NF. This loss function imposes a physics-informed data fidelity that includes both the forward model in $\mathbf{H}_{i}$ as well as a Fourier-noise-specific $|\mathbf{k}|$ weighting as in [5]. We use TV loss to regularize the conventional pixel grid representation, while for the NF representation, we rely instead on the network’s implicit regularization and set $\lambda=0$. Training is performed only on the SOS representation parameters $\phi$ but requiring accurate aberration reconstruction across all delays and image patches leads to sufficiently accurate PSFs that the final multi-channel deconvolution produces high-quality recovered IP image patches $x_{i}$. Finally, these patches are merged into a single image by applying a Gaussian window with a FWHM of 1.5 mm [5] for smooth interpatch transitions and adding the patches back to the IP image at corresponding patch locations. We use a delay count of $M=16$ by default in this paper (see ablations in Sec. 4.4). See Sec. S1.5 in the supplement for details on our image patches and stitching.
Our method is implemented with PyTorch [21] and available on GitHub (see project page). We trained for 10 epochs for NF and 30 epochs for PG with Adam optimizer [16] and learning rates of $5\times 10^{-3}$ and $1\times 10^{-1}$ respectively on an NVIDIA RTX A6000 GPU.
@@ 4.1 Numerical Simulations
To quantitatively benchmark the performance of our method with ground truth references, we designed a set of numerical phantoms with 5 different SOS maps (see Sec. S1.1 for details), one of which is shown in Fig. 4a for visualization.
@@
The wavefront errors modeled/reconstructed (sample location marked by white “x" on SOS in Fig. 4a) by different methods are shown in Fig. 4b. To use the terminology of [5], the wavefront errors can be decomposed into a Fourier series, where the components degrade the IP image differently. Conventional DAS assumes a uniform SOS thus a constant wavefront (red), while Dual-SOS DAS assumes an additional body SOS and is able to capture the zeroth and first-order term accurately (yellow). APACT (green) includes only the zeroth and second order Fourier series terms in its wavefront model (see [5] and Sec. S2 in the supplement for details), missing both the first order (the cosine component captured by dual-SOS in yellow) as well as higher order terms capturing finer detail. In comparison, our methods can accurately model lower and higher-order components in the wavefront (purple and blue), compared with the truth wavefront (black).
@@
These wavefront errors are directly reflected in image reconstruction quality in Fig. 4c. Conventional DAS, which assumes a uniform SOS of 1510.0 m/s, not only suffers from doubling artifacts (see insets) due to the ignorance of higher order terms but also causes the image to shrink (see vertical shift of the circle and boundary in green box). This shrinking is caused by the missing first-order term in the wavefront, which leads to off-center PSFs that shift the image patches toward the center of the tissue. Note also that for real samples this uniform SOS value is tuned by generating many reconstructions and requiring the user to select the best image, significantly increasing the reconstruction time in practice. Dual-SOS DAS assumes a body SOS of 1561.0 m/s (also hand-tuned with search time not included in reported time) and is able to fix the shrinking problem but artifacts remain. The image reconstructed by APACT is also shrunk (see green box and the image quality metrics) due to ignorance of the first-order term. APACT oversmooths the SOS and loses key features in the SOS, e.g., the orange liver-like region and the hole in the upper right, and is also slow and expensive due to the exhaustive search for wavefronts coefficients. Our method reconstructs a more accurate SOS, corresponding to a significant improvement in image quality over traditional methods. Image patches show fewer aberrations than all previous methods, reflected in high PSNR and SSIM scores. Our method with NF takes only 40 seconds, compared to 23 minutes for APACT. Our method with PG produces decent IP images in similar compute time but suffers from stripe-like SOS artifacts that originate from the path-integral nature of the wavefront calculation (see Eq. 3) in the straight acoustic ray model. We also show the image reconstructed by multi-channel deconvolution using the ground truth SOS (SOS oracle) as an upper limit baseline for image quality and speed of USCT-enhanced reconstruction. See Tab. 1 for quantitative evaluation of all the methods, averaged over five numerical phantoms. Supplementary videos show the convergence of the IP image and SOS for each experiment.